OpenStudy (anonymous):

Please Help!?!? Sketch the graph of f(x)={ x^2+4x+2, x<-2 {1-4x-x^2, x>_ -2 (It is one big bracket) also if you can (a) Is f continuous at x=-2? (b) Is f differentiable at x=-2? Explain

OpenStudy (anonymous):

Please help me, i really have no idea how to do this

OpenStudy (phi):

pick some numbers for x, and figure out what f(x) for example, at x=0, f(x) = 0^2 + 4*0 + 2 = 0+0+2 = 2 so (0,2) is a point on this parabola. try x=-1, x=-2 x= -3, note that they only want the curve for x < -2, but it does not hurt to plot the whole thing and then "erase" the part they don't want.

OpenStudy (anonymous):

how do i graph it though? like what should the graph look like?

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

is it the same for the type of function I have to graph?

OpenStudy (phi):

yes

OpenStudy (phi):

the example is closer to your second equation 1-4x-x^2 (this one looks like a "frown"

OpenStudy (anonymous):

lol its just i didn't know how to express the inequality

OpenStudy (anonymous):

because it is a greater than sign

OpenStudy (phi):

f(x)={ x^2+4x+2, x<-2 {, x≥ -2 this means that when x is less than -2, you use the top equation when x is -2 or bigger, you use the bottom equation For the time being, ignore the x<-2 or x≥ -2 just plot the top equation (use a table of values like in the video) then do the same for the bottom equation after you are done, at x<-2, erase the bottom equation's curve at at x≥ -2 , erase the top equation's curve. It will look something like this